There are two goods X and Y. Tabithas Utilities are given by U = 1/3X2 + 2 Y1/2. If the price of X is $2 and the Price of Y is $0.75 What is Tabithas optimum bundle?
1
Expert's answer
2020-10-07T07:12:03-0400

"U= 1\/3 x^2 + y^{1\/2}"

 and 

x=2, y=0.75=3/4

"MR_{s} = \frac {MU_{x}}{MU_{y}}
 = \frac {2/3x}{\frac{2}{2}y^{1/2}}
 =\frac {2/3x}{y^{1/2}}"

="\frac {2}{3} xy^2"

MRs is not Diminishing

if x is consumed then,


  • if x is consumed then


"x'=\frac{m}{p_{x}}" x` "=" "\frac {m}{P_{x}}""=" "\frac {2}{p_{x}}"

"U= 1/3(\frac{m}{2})^2 = \frac {m^2}{12}"


  • if y is consumer then,

"y` =\frac {m}{p_{y}} = \frac {m}{0.75} = \frac {4}{3} m"

"U= 2 (\frac {4}{3}m)^{1/2} = 4 \sqrt (\frac{m}{3})"